• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name khandaker

    Log entry time 19:15:31 on February28,2002

    Entry number 626

    keyword=Q1 quenched

    18:09 HRS-L Q1 quenched. The liquid He flow was reading -1.77 l/m and could not reset the magnet. Tech on-call said that it takes several minutes after a quench before the proper flow rate (~100 l/m) gets stabilized.
    After Q1 reset, noticed that Q1 current was fluctuating by deltaI~0.5 A. This was triggering off the HV alarm every 10-15 sec. Discussion w/ the on-call expert,
    this amount of fluctuation corresponds to 4x10^-4 GeV/c variations. This is O.K.
    for now.